### v1

@@ 2.1 Weak formulation and preliminary properties
$g_{1}\leq g(s)\leq g_{2}\qquad\forall s\in\mathbb{R}^{+},$ (5)
@@
On the permeable sub-boundary $\Gamma_{\mathrm{m}}$ we will define the following quantity
$\lambda:=-(\boldsymbol{\sigma}\boldsymbol{n})\cdot\boldsymbol{n},$ (6)
which will be treated as a Lagrange multiplier. We then proceed to define the following functional spaces for fluid velocity, pressure, the Lagrange multiplier, and concentration, respectively {linenomath}
@@
$\boldsymbol{\mathrm{H}}_{0}:=\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)\;:\;\boldsymbol{v}\cdot\boldsymbol{t}=0\text{ on }\Gamma_{\mathrm{m}},\,\boldsymbol{v}=\boldsymbol{0}\text{ on }\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{wall}}\},\qquad Q:=L^{2}(\Omega),\qquad W:=\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}}),$
@@
$\int_{\Omega}\mu_{0}\nabla\boldsymbol{u}:\nabla\boldsymbol{v}-\int_{\Omega}p\nabla\cdot\boldsymbol{v}+\int_{\Omega}\rho_{0}\left(\boldsymbol{u}\cdot\nabla\boldsymbol{u}\right)\cdot\boldsymbol{v}+\langle\lambda,\boldsymbol{v}\cdot\boldsymbol{n}\rangle_{\Gamma_{\mathrm{m}}}=0.$ (7)
@@
$\nabla\cdot\left(\theta\boldsymbol{u}-D_{0}\nabla\theta\right)=0\quad\text{in}\;\Omega.$ (8)
Then, testing (8) against $\tau\in Z_{0}$, integrating by parts and using the boundary conditions (2b) and (2e) we obtain
$D_{0}\int_{\Omega}\nabla\theta\cdot\nabla\tau-\int_{\Omega}\theta(\boldsymbol{u}\cdot\nabla\tau)+(\theta(\boldsymbol{u}\cdot\boldsymbol{n}),\tau)_{\Gamma_{\mathrm{out}}}=0\quad\forall\tau\in Z_{0}.$
@@
$\langle\xi,\boldsymbol{u}\cdot\boldsymbol{n}\rangle_{\Gamma_{\mathrm{m}}}=\langle\xi,g(\theta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall\xi\in W,$ (9)
whereas the zero tangential velocity condition (2f) is imposed strongly on the velocity space.
Let us remark that for $\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)$, we have that $\boldsymbol{v}\cdot\boldsymbol{n}\in\mathrm{H}^{1/2}(\Gamma_{j})$, for $\Gamma_{j}\subset\partial\Omega$. But if $\boldsymbol{v}\cdot\boldsymbol{n}$ vanishes on $O(\Gamma_{j})$ (an open subset of $\partial\Omega$ containing $\overline{\Gamma_{i}}$), then $\boldsymbol{v}\cdot\boldsymbol{n}\in\mathrm{H}^{1/2}_{00}(\Gamma_{j})$ (see, e.g., [21, Sect. 1.5]). This means that if the boundary configuration is such that $\Gamma_{\mathrm{m}}$ is surrounded by $\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{wall}}$ (which is not the case of Figure 2.1 since $\Gamma_{\mathrm{m}}$ is adjacent with $\Gamma_{\mathrm{out}}$), then we would have $\boldsymbol{v}\cdot\boldsymbol{n}\in\mathrm{H}^{1/2}_{00}(\Gamma_{\mathrm{m}})$ and we would need $\lambda\in\mathrm{H}^{-1/2}_{00}(\Gamma_{\mathrm{m}})$.
Furthermore, we introduce the following bilinear and trilinear forms {linenomath}
$a(\boldsymbol{u},\boldsymbol{v}):=\mu_{0}\int_{\Omega}\nabla\boldsymbol{u}:\nabla\boldsymbol{v},\qquad\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v}):=\rho_{0}\int_{\Omega}(\boldsymbol{w}\cdot\nabla\boldsymbol{u})\cdot\boldsymbol{v},$
$b(\boldsymbol{v},(q,\xi)):=-\int_{\Omega}q\nabla\cdot\boldsymbol{v}+\langle\xi,\boldsymbol{v}\cdot\boldsymbol{n}\rangle_{\Gamma_{\mathrm{m}}},$
$c(\theta,\tau):=D_{0}\int_{\Omega}\nabla\theta\cdot\nabla\tau,\qquad\widetilde{c}(\boldsymbol{w};\theta,\tau):=-\int_{\Omega}\theta(\boldsymbol{w}\cdot\nabla\tau)+(\theta(\boldsymbol{w}\cdot\boldsymbol{n}),\tau)_{\Gamma_{\mathrm{out}}}.$
@@
$\boldsymbol{U}=\boldsymbol{u}_{\mathrm{in}}\quad\text{on $\Gamma_{\mathrm{in}}$},\qquad\boldsymbol{U}=\boldsymbol{0}\quad\text{on $\Gamma_{\mathrm{wall}}$},\qquad\mathop{\mathrm{div}}\nolimits\,\boldsymbol{U}=0\quad\text{in $\Omega$},\qquad\Theta=\theta_{\mathrm{in}}\quad\text{on $\Gamma_{\mathrm{in}}$}.$ (10)
@@
$a(\boldsymbol{U}+\boldsymbol{u}_{0},\boldsymbol{v})+\widetilde{a}(\boldsymbol{U}+\boldsymbol{u}_{0};\boldsymbol{U}+\boldsymbol{u}_{0},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=0$ $\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$ (11)
@@
Next we stress that the bilinear and trilinear forms considered in the above weak formulation are uniformly bounded. It suffices to apply Hölder’s inequality, Sobolev embeddings and trace inequalities (see, for instance, [39, Section 1.1] or [6, Section 9.2]): {linenomath}
$|a(\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{u}\|_{1,\Omega}\|\boldsymbol{v}\|_{1,\Omega}$ $\boldsymbol{u},\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),$ (12a)
$|b(\boldsymbol{v},(q,\xi))|$ $\lesssim\|\boldsymbol{v}\|_{1,\Omega}(\|q\|_{0,\Omega}+\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}})$ $\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),q\in L^{2}(\Omega),\xi\in\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}}),$ (12b)
$|c(\psi,\tau)|$ $\lesssim\|\psi\|_{1,\Omega}\|\tau\|_{1,\Omega}$ $\psi,\tau\in\mathrm{H}^{1}(\Omega),$ (12c)
$|\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{w}\|_{1,\Omega}\,\|\boldsymbol{u}\|_{1,\Omega}\,\|\boldsymbol{v}\|_{1,\Omega},$ $\boldsymbol{w},\boldsymbol{u},\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)$ (12d)
$|\widetilde{c}(\boldsymbol{w};\theta,\tau)|$ $\lesssim\|\boldsymbol{w}\|_{1,\Omega}\,\|\tau\|_{1,\Omega}\,\,\|\theta\|_{1,\Omega},$ $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),\theta,\tau\in\mathrm{H}^{1}(\Omega).$ (12e)
In particular, for (12e) we have used that {linenomath}
$\left|\int_{\Gamma_{\mathrm{out}}}\theta(\boldsymbol{w}\cdot\boldsymbol{n})\tau\right|$ $\leq\|\tau\|_{L^{4}(\Gamma_{\mathrm{out}})}\,\|\boldsymbol{w}\|_{0,\Gamma_{\mathrm{out}}}\,\|\theta\|_{L^{4}(\Gamma_{\mathrm{out}})}$
$\leq\|\tau\|_{L^{4}(\partial\Omega)}\,\|\boldsymbol{w}\|_{0,\partial\Omega}\,\|\theta\|_{L^{4}(\partial\Omega)}.$
We also note that thanks to the vector and scalar forms of Poincaré inequality, we have the ellipticity for the bilinear forms $a(\cdot,\cdot)$ and $c(\cdot,\cdot)$: {linenomath}
$|a(\boldsymbol{v},\boldsymbol{v})|$ $\gtrsim\|\boldsymbol{v}\|_{1,\Omega}^{2}$ $\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),$ (13a)
$|c(\tau,\tau)|$ $\gtrsim\|\tau\|_{1,\Omega}^{2}$ $\forall\tau\in\mathrm{H}^{1}(\Omega),$ (13b)
respectively.
@@
$\boldsymbol{\mathrm{X}}^{g}$ :=$:=\left\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}:b(\boldsymbol{v},(q,\xi))=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\ \forall(q,\xi)\in Q\times W\right\}$
=$=\left\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}:\nabla\cdot\boldsymbol{v}=0\text{ in }\Omega,\quad\boldsymbol{v}\cdot\boldsymbol{n}=g(\zeta)\text{ on }\Gamma_{\mathrm{m}}\right\},$ (14)
=$=\left\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega):\nabla\cdot\boldsymbol{v}=0\text{ in }\Omega,\quad\boldsymbol{v}=\boldsymbol{u}_{\mathrm{in}}\text{ on }\Gamma_{\mathrm{in}},\quad\boldsymbol{v}\cdot\boldsymbol{n}=g(\zeta)\text{ on }\Gamma_{\mathrm{m}},\quad\boldsymbol{v}=\boldsymbol{0}\text{ on }\Gamma_{\mathrm{wall}}\right\}.$
@@
$\widetilde{c}(\boldsymbol{w};\tau,\tau)=\frac{1}{2}[(\boldsymbol{w}\cdot\boldsymbol{n},\tau^{2})_{\Gamma_{\mathrm{out}}}-(\boldsymbol{w}\cdot\boldsymbol{n},\tau^{2})_{\Gamma_{\mathrm{m}}}]\qquad\forall\boldsymbol{w}\in\boldsymbol{\mathrm{X}}^{g},\;\forall\tau\in Z_{0},$ (15)
@@
$\widetilde{c}(\boldsymbol{w};\tau,\tau)\geq 0,\qquad\boldsymbol{w}\in\boldsymbol{\mathrm{X}}^{g},\tau\in Z_{0}.$ (16)
@@ 3 Well-posedness of the continuous problem
For the analysis of existence and uniqueness of solution we will use a fixed-point argument separating the solution between two uncoupled problems. First consider the Navier–Stokes equations, which consist in finding, for a given $\zeta=\zeta_{0}+\Theta$ (with $\zeta_{0}\in Z_{0}$), the tuple $(\boldsymbol{u}_{0}+\boldsymbol{U},p,\lambda)\in\boldsymbol{\mathrm{H}}\times Q\times W$ such that {linenomath}
$a(\boldsymbol{u},\boldsymbol{v})+\widetilde{a}(\boldsymbol{u};\boldsymbol{u},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=0\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$ (17a)
$b(\boldsymbol{u},(q,\xi))$ $=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$ (17b)
@@
$c(\theta_{0},\tau)+\widetilde{c}(\boldsymbol{w};\theta_{0},\tau)=-c(\Theta,\tau)-\widetilde{c}(\boldsymbol{w};\Theta,\tau)\quad\forall\tau\in Z_{0}.$ (18)
@@ 3.1 Well-posedness of the Navier–Stokes equations with membrane boundary conditions
In order to address the unique solvability of (17a)-(17b), we use a linear Stokes problem with membrane boundary conditions, the Banach fixed-point theorem, and the Babuška–Brezzi theory for saddle-point problems [18]. For this we follow the analysis from [13]. Let us consider the problem of finding, for a given $\zeta\in Z$ and a given $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}_{0}$, the tuple $(\boldsymbol{u}_{0},p,\lambda)\in\boldsymbol{\mathrm{H}}_{0}\times Q\times W$ such that {linenomath}
$a(\boldsymbol{u}_{0},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=-a(\boldsymbol{U},\boldsymbol{v})-\widetilde{a}(\boldsymbol{w}+\boldsymbol{U};\boldsymbol{w}+\boldsymbol{U},\boldsymbol{v})\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$ (19a)
$b(\boldsymbol{u}_{0},(q,\xi))$ $=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$ (19b)
In order to show that this linear Stokes system is well-posed we follow arguments similar to [41, Lemma 3.1], [15, 30], and [17, Section 2.4.3]. We start with the following result (the proof is carried out in a standard way, but we present it for sake of completeness).
@@
Thanks to the Riesz representation theorem, for a given $\xi\in W$ there exists $\tilde{\xi}\in\mathrm{H}^{1/2}(\Gamma_{\mathrm{m}})$ such that $\|\tilde{\xi}\|_{1/2,\Gamma_{\mathrm{m}}}=\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}}$. For a given pair $(q,\xi)\in Q\times W$, let us consider the following auxiliary Stokes problem with mixed boundary conditions {linenomath}
[…]
$(\boldsymbol{\nabla}\hat{\boldsymbol{v}}-\zeta\boldsymbol{\mathrm{I}})\boldsymbol{n}$ $=\boldsymbol{0}\quad\text{on }\Gamma_{\mathrm{out}},$ (20)
@@
Thanks to [19], we can assert that there exists a unique velocity solution to (20), for which there holds
$\nabla\cdot\hat{\boldsymbol{v}}=q,\quad(\hat{\boldsymbol{v}}\cdot\boldsymbol{n})|_{\Gamma_{\mathrm{m}}}=\tilde{\xi},\quad(\hat{\boldsymbol{v}}\cdot\boldsymbol{t})|_{\Gamma_{\mathrm{m}}}=0,$ (21a)
$\|\hat{\boldsymbol{v}}\|_{1,\Omega}\lesssim\|q\|_{0,\Omega}+\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}},$ (21b)
and so $\hat{\boldsymbol{v}}\in\boldsymbol{\mathrm{H}}_{0}\setminus\{\boldsymbol{0}\}$. In this way we can write

### v2

@@ 2.1 Weak formulation and preliminary properties
(5) $g_{1}\leq g(s)\leq g_{2}\qquad\forall s\in\mathbb{R}^{+},$
@@
On the permeable sub-boundary $\Gamma_{\mathrm{m}}$ we define the following quantity
(6) $\lambda:=-(\boldsymbol{\sigma}\boldsymbol{n})\cdot\boldsymbol{n},$
which is treated as a Lagrange multiplier. We then proceed to define the following functional spaces for fluid velocity, pressure, the Lagrange multiplier, and concentration, respectively
@@
$\boldsymbol{\mathrm{H}}_{0}:=\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)\;:\;\boldsymbol{v}\cdot\boldsymbol{t}=0\text{ on }\Gamma_{\mathrm{m}},\,\boldsymbol{v}=\boldsymbol{0}\text{ on }\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{wall}}\},\qquad Q:=\mathrm{L}^{2}(\Omega),\qquad W:=\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}}),$
@@
(7) $\int_{\Omega}\mu_{0}\nabla\boldsymbol{u}:\nabla\boldsymbol{v}-\int_{\Omega}p\nabla\cdot\boldsymbol{v}+\int_{\Omega}\rho_{0}\left(\boldsymbol{u}\cdot\nabla\boldsymbol{u}\right)\cdot\boldsymbol{v}+\langle\lambda,\boldsymbol{v}\cdot\boldsymbol{n}\rangle_{\Gamma_{\mathrm{m}}}=0.$
@@
(8) $\nabla\cdot\left(\theta\boldsymbol{u}-D_{0}\nabla\theta\right)=0\quad\text{in}\;\Omega.$
Then, testing (8) against $\tau\in Z_{0}$, integrating by parts and using (2b) and (2e) we obtain
$D_{0}\int_{\Omega}\nabla\theta\cdot\nabla\tau-\int_{\Omega}\theta(\boldsymbol{u}\cdot\nabla\tau)+\langle\theta(\boldsymbol{u}\cdot\boldsymbol{n}),\tau\rangle_{\Gamma_{\mathrm{out}}}=0\quad\forall\tau\in Z_{0},$
where the well-definedness of the last term on the left-hand side is addressed later in (12e).
@@
(9) $\langle\xi,\boldsymbol{u}\cdot\boldsymbol{n}\rangle_{\Gamma_{\mathrm{m}}}=\langle\xi,g(\theta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall\xi\in W,$
whereas the zero tangential velocity condition (2f) is imposed strongly on the velocity space. Furthermore, we introduce the following bilinear and trilinear forms
$a(\boldsymbol{u},\boldsymbol{v}):=\mu_{0}\int_{\Omega}\nabla\boldsymbol{u}:\nabla\boldsymbol{v},\qquad\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v}):=\rho_{0}\int_{\Omega}(\boldsymbol{w}\cdot\nabla\boldsymbol{u})\cdot\boldsymbol{v},\qquad c(\theta,\tau):=D_{0}\int_{\Omega}\nabla\theta\cdot\nabla\tau,$
$b(\boldsymbol{v},(q,\xi)):=-\int_{\Omega}q\nabla\cdot\boldsymbol{v}+\langle\xi,\boldsymbol{v}\cdot\boldsymbol{n}\rangle_{\Gamma_{\mathrm{m}}},\qquad\widetilde{c}(\boldsymbol{w};\theta,\tau):=-\int_{\Omega}\theta(\boldsymbol{w}\cdot\nabla\tau)+\langle\theta(\boldsymbol{w}\cdot\boldsymbol{n}),\tau\rangle_{\Gamma_{\mathrm{out}}}.$
@@
(10) $\boldsymbol{U}=\boldsymbol{u}_{\mathrm{in}}\;\text{on}\;\Gamma_{\mathrm{in}},\quad\boldsymbol{U}=\boldsymbol{0}\;\text{on}\;\Gamma_{\mathrm{wall}},\quad\boldsymbol{U}\cdot\boldsymbol{t}=0\;\text{on}\;\Gamma_{\mathrm{m}},\quad\mathop{\mathrm{div}}\nolimits\,\boldsymbol{U}=0\;\text{in}\;\Omega,\quad\Theta=\theta_{\mathrm{in}}\;\text{on}\;\Gamma_{\mathrm{in}}.$
@@
(11) $a(\boldsymbol{U}+\boldsymbol{u}_{0},\boldsymbol{v})+\widetilde{a}(\boldsymbol{U}+\boldsymbol{u}_{0};\boldsymbol{U}+\boldsymbol{u}_{0},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=0$ $\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$
@@
Next we stress that the bilinear and trilinear forms considered in the above weak formulation are uniformly bounded. It suffices to apply Hölder’s inequality, Sobolev embeddings and trace inequalities (see, for instance, [36, Section 1.1]):
(12a) $|a(\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{u}\|_{1,\Omega}\|\boldsymbol{v}\|_{1,\Omega}$ $\boldsymbol{u},\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),$
(12b) $|b(\boldsymbol{v},(q,\xi))|$ $\lesssim\|\boldsymbol{v}\|_{1,\Omega}(\|q\|_{0,\Omega}+\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}})$ $\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),q\in L^{2}(\Omega),\xi\in\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}}),$
(12c) $|c(\psi,\tau)|$ $\lesssim\|\psi\|_{1,\Omega}\|\tau\|_{1,\Omega}$ $\psi,\tau\in\mathrm{H}^{1}(\Omega),$
(12d) $|\widetilde{a}(\boldsymbol{w};\boldsymbol{u},\boldsymbol{v})|$ $\lesssim\|\boldsymbol{w}\|_{1,\Omega}\,\|\boldsymbol{u}\|_{1,\Omega}\,\|\boldsymbol{v}\|_{1,\Omega},$ $\boldsymbol{w},\boldsymbol{u},\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega)$
(12e) $|\widetilde{c}(\boldsymbol{w};\theta,\tau)|$ $\lesssim\|\boldsymbol{w}\|_{1,\Omega}\,\|\tau\|_{1,\Omega}\,\,\|\theta\|_{1,\Omega},$ $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),\theta,\tau\in\mathrm{H}^{1}(\Omega).$
In particular, for (12e) we have used that
$\left|\int_{\Gamma_{\mathrm{out}}}\theta(\boldsymbol{w}\cdot\boldsymbol{n})\tau\right|\leq\|\tau\|_{L^{4}(\Gamma_{\mathrm{out}})}\,\|\boldsymbol{w}\|_{0,\Gamma_{\mathrm{out}}}\,\|\theta\|_{L^{4}(\Gamma_{\mathrm{out}})}\leq\|\tau\|_{L^{4}(\partial\Omega)}\,\|\boldsymbol{w}\|_{0,\partial\Omega}\,\|\theta\|_{L^{4}(\partial\Omega)}.$
We also note that thanks to the vector and scalar forms of Poincaré inequality, we have, respectively, the ellipticity for the bilinear forms $a(\cdot,\cdot)$ and $c(\cdot,\cdot)$ as
(13a) $|a(\boldsymbol{v},\boldsymbol{v})|$ $\gtrsim\|\boldsymbol{v}\|_{1,\Omega}^{2}$ $\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}^{1}(\Omega),$
(13b) $|c(\tau,\tau)|$ $\gtrsim\|\tau\|_{1,\Omega}^{2}$ $\forall\tau\in\mathrm{H}^{1}(\Omega).$
@@
$\boldsymbol{\mathrm{X}}^{g}$ $:=\left\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}:b(\boldsymbol{v},(q,\xi))=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\ \forall(q,\xi)\in Q\times W\right\}$
(14) $=\left\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}:\nabla\cdot\boldsymbol{v}=0\text{ in }\Omega,\quad\boldsymbol{v}\cdot\boldsymbol{n}=g(\zeta)\text{ on }\Gamma_{\mathrm{m}}\right\},$
[…]
@@
(15) $\widetilde{c}(\boldsymbol{w};\tau,\tau)=\frac{1}{2}[(\boldsymbol{w}\cdot\boldsymbol{n},\tau^{2})_{\Gamma_{\mathrm{out}}}-(\boldsymbol{w}\cdot\boldsymbol{n},\tau^{2})_{\Gamma_{\mathrm{m}}}]\qquad\forall\boldsymbol{w}\in\boldsymbol{\mathrm{X}}^{g},\;\forall\tau\in Z_{0},$
@@
(16) $\widetilde{c}(\boldsymbol{w};\tau,\tau)\geq 0,\qquad\boldsymbol{w}\in\boldsymbol{\mathrm{X}}^{g},\tau\in Z_{0}.$
@@ 3 Well-posedness of the continuous problem
For the analysis of existence and uniqueness of solution we use a fixed-point argument separating the solution between two uncoupled problems. First consider the Navier–Stokes equations, which consist in finding, for a given $\zeta=\zeta_{0}+\Theta$ (with $\zeta_{0}\in Z_{0}$), the tuple $(\boldsymbol{u}_{0}+\boldsymbol{U},p,\lambda)\in\boldsymbol{\mathrm{H}}\times Q\times W$ such that
(17a) $a(\boldsymbol{u},\boldsymbol{v})+\widetilde{a}(\boldsymbol{u};\boldsymbol{u},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=0\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$
(17b) $b(\boldsymbol{u},(q,\xi))$ $=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$
@@
(18) $c(\theta_{0},\tau)+\widetilde{c}(\boldsymbol{w};\theta_{0},\tau)=-c(\Theta,\tau)-\widetilde{c}(\boldsymbol{w};\Theta,\tau)\quad\forall\tau\in Z_{0}.$
@@ 3.1 Well-posedness of the Navier–Stokes equations with membrane boundary conditions
In order to address the unique solvability of (17a)-(17b), we use a linear Stokes problem with membrane boundary conditions, the Banach fixed-point theorem, and the Babuška–Brezzi theory for saddle-point problems. For this we follow the analysis from [12]. Let us consider the problem of finding, for a given $\zeta\in Z$ and a given $\boldsymbol{w}\in\boldsymbol{\mathrm{H}}_{0}$, the tuple $(\boldsymbol{u}_{0},p,\lambda)\in\boldsymbol{\mathrm{H}}_{0}\times Q\times W$ such that
(19a) $a(\boldsymbol{u}_{0},\boldsymbol{v})+b(\boldsymbol{v},(p,\lambda))$ $=-a(\boldsymbol{U},\boldsymbol{v})-\widetilde{a}(\boldsymbol{w}+\boldsymbol{U};\boldsymbol{w}+\boldsymbol{U},\boldsymbol{v})\quad\forall\boldsymbol{v}\in\boldsymbol{\mathrm{H}}_{0},$
(19b) $b(\boldsymbol{u}_{0},(q,\xi))$ $=\langle\xi,g(\zeta)\rangle_{\Gamma_{\mathrm{m}}}\quad\forall(q,\xi)\in Q\times W.$
In order to show that this linear Stokes system is well-posed we follow arguments similar to [38, Lemma 3.1], [15, 27]. We start with the following result (the proof is carried out in a standard way, but we present it for sake of completeness).
@@
Thanks to the Riesz representation theorem, for a given $\xi\in W$ there exists $\tilde{\xi}\in\mathrm{H}^{1/2}(\Gamma_{\mathrm{m}})$ such that $\|\tilde{\xi}\|_{1/2,\Gamma_{\mathrm{m}}}=\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}}$. For a given pair $(q,\xi)\in Q\times W$, let us consider the following auxiliary Stokes problem with mixed boundary conditions
$-\boldsymbol{\Delta}\hat{\boldsymbol{v}}+\nabla\zeta=\boldsymbol{0}\quad\text{and}\quad\nabla\cdot\hat{\boldsymbol{v}}$ $=q\quad\text{in }\Omega,$
(20) $(\boldsymbol{\nabla}\hat{\boldsymbol{v}}-\zeta\boldsymbol{\mathrm{I}})\boldsymbol{n}$ $=\boldsymbol{0}\quad\text{on }\Gamma_{\mathrm{out}},$
@@
For the compatibility of the Dirichlet data in (20), we can proceed similarly as in [30, 18] (see also [28]) and use a cut-off and mollification argument. Let us first define an extension of $\tilde{\xi}$ on $\Gamma_{*}:=\Gamma_{\mathrm{m}}\cup\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{wall}}$ as follows
$\tilde{\xi}_{0}(\boldsymbol{x}):=\begin{cases}\tilde{\xi}(\boldsymbol{x})&\boldsymbol{x}\in\Gamma_{\mathrm{m}},\\ 0&\boldsymbol{x}\in\Gamma_{*}\setminus\Gamma_{\mathrm{m}},\end{cases}$
and note that the extension is linear in $\tilde{\xi}$. Also, $\tilde{\xi}_{0}\in\mathrm{L}^{2}(\Gamma_{*})$, but it is not necessarily in $\mathrm{H}^{1/2}(\Gamma_{*})$ because of the jump across the left-bottom corner belonging to $\partial\Gamma_{\mathrm{m}}$. Next we consider a smooth cut-off function $\eta_{\epsilon}\in C^{\infty}(\Gamma_{*})$ satisfying $0\leq\eta_{\epsilon}\leq 1$, $\eta_{\epsilon}\equiv 1$ on the interior of $\Gamma_{\mathrm{m}}$, and $\text{supp}(\eta_{\epsilon})\subset\Gamma_{*}$. Its transition from 1 to 0 happens in a small layer of width $\epsilon$ near the aforementioned corner in $\partial\Gamma_{\mathrm{m}}\subset\Gamma_{*}$. In addition, we denote by $\rho_{\epsilon}$ a standard smooth mollifier defined on local charts of $\Gamma_{*}$, and define the function
$\xi_{\epsilon}^{*}:=\eta_{\epsilon}(\tilde{\xi}_{0}*\rho_{\epsilon}),$
where the convolution smooths out the jump across the left-bottom corner in $\partial\Gamma_{\mathrm{m}}$. Then we have $\xi_{\epsilon}^{*}\in\mathrm{H}^{1/2}(\Gamma_{*})$, $\text{supp}(\xi_{\epsilon}^{*})\subset\Gamma_{*}$, $\xi_{\epsilon}^{*}\equiv\tilde{\xi}$ on the interior of $\Gamma_{\mathrm{m}}$ (away from the $\epsilon$-transition layer), and $\xi_{\epsilon}^{*}$ vanishes near $\partial\Gamma_{*}$. Therefore $\xi_{\epsilon}^{*}\in\mathrm{H}^{1/2}_{00}(\Gamma_{*})$.
We further notice that the map $\tilde{\xi}\mapsto\xi_{\epsilon}^{*}$ is linear and bounded
(21) $\|\xi_{\epsilon}^{*}\|_{1/2,00;\Gamma_{*}}\leq C_{\epsilon}\|\tilde{\xi}\|_{1/2,\Gamma_{\mathrm{m}}}=C_{\epsilon}\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}},$
with $C_{\epsilon}>0$ a constant depending only on $\epsilon$ and on the geometry of $\Gamma_{\mathrm{m}}$ and $\Gamma_{*}$.
Then, the boundary conditions in (20) are rewritten in a well-defined manner as
(22) $(\boldsymbol{\nabla}\hat{\boldsymbol{v}}-\zeta\boldsymbol{\mathrm{I}})\boldsymbol{n}=\boldsymbol{0}\quad\text{on }\Gamma_{\mathrm{out}},\qquad\hat{\boldsymbol{v}}=\xi_{\epsilon}^{*}\boldsymbol{n}\quad\text{on }\Gamma_{*}=\partial\Omega\setminus\Gamma_{\mathrm{out}},$
with $\xi_{\epsilon}^{*}\boldsymbol{n}\in\boldsymbol{\mathrm{H}}^{1/2}_{00}(\Gamma_{*})$, and thanks to [17], we can assert that there exists a unique velocity solution to (20) (with mixed boundary conditions as in (22)), for which it is straightforward to verify that there holds
(23a) $\nabla\cdot\hat{\boldsymbol{v}}=q,\quad(\hat{\boldsymbol{v}}\cdot\boldsymbol{n})|_{\Gamma_{\mathrm{m}}}=\tilde{\xi},\quad(\hat{\boldsymbol{v}}\cdot\boldsymbol{t})|_{\Gamma_{\mathrm{m}}}=0,$
(23b) $\|\hat{\boldsymbol{v}}\|_{1,\Omega}\lesssim\|q\|_{0,\Omega}+\|\xi\|_{-1/2,\Gamma_{\mathrm{m}}},$
(where we have also used (21)). In this way we have that $\hat{\boldsymbol{v}}\in\boldsymbol{\mathrm{H}}_{0}\setminus\{\boldsymbol{0}\}$, and thus we can write